OpenStudy (girondasearch):

could someone help me with this one any help is appreciated thank let f (x)=x^2 + 6 and g (x) =x+8/x find (g o f) (-7) a) 384 /7 b) -57/7 c) 295 /49 d) 63/55 its not C i got that one wrong

OpenStudy (phi):

(g o f) (-7) means first figure out f(-7) f (x)=x^2 + 6 f(-7) = (-7)^2 + 6 what do you get if you simplify that ?

OpenStudy (phi):

do you know how to figure out -7* -7 + 6 ?

OpenStudy (girondasearch):

sorry it's been a long time since i have done calculus i'm getting back into pre calculus that is where they are starting me so i'm trying to refresh my mind

OpenStudy (phi):

the arithmetic follows rules PEDMAS order of operations are done in the order: parens exponents (powers) division/multiply add/subtract in this problem (-7)^2 + 6 you do the the -7*-7 first then you add 6

OpenStudy (phi):

once you get a number, you "use" that number in place of x in g(x) = (x+8)/x

OpenStudy (girondasearch):

ok thank you this helps me i don't know why i was road blocking myself big thank you

likeabossssssss (likeabossssssss):

welcome to os

OpenStudy (girondasearch):

thank you very much

likeabossssssss (likeabossssssss):

np